# Constraint-Generation Policy Optimization (CGPO): Nonlinear Programming for Policy Optimization in Mixed Discrete-Continuous MDPs

### v1

@@ Introduction
Table 1: Problem classes of the inner/outer optimization problems in Algorithm 1 for different expressivity classes of dynamics/reward (columns) and policies (rows).
$reorder=4-stock$
$F=0.6-15.4\theta-2.3\omega+100\theta^{2}-1.5\omega^{2}$
Figure 2: Optimal S/PWS-S policy for inventory (left), PWS-C policy for reservoir (middle), Q policy for VTOL (right).
Implementation Details
Relational Dynamic influence Diagram Language (RDDL) is a structured planning domain description language particularly well-suited for modelling problems with stochastic effects (Sanner 2010). Our domains are therefore coded in RDDL (please see supplementary material for the domain descriptions) and we wrote a general-purpose routine for compiling RDDL code into a Gurobi mixed integer program formulation using the pyRDDLGym interface (Taitler et al. 2022). In order to compute the tightest possible bounds on decision variables in the MIP compilation, we use interval arithmetic (Hickey, Ju, and Van Emden 2001).
Empirical Evaluation
We now empirically validate the soundness and performance of Algorithm 1 on several standard MDPs. The goal of the empirical evaluation is to answer the following research questions:
  1. Q1:
@@
  2. Q2:
    How do different policy expressivity classes perform per constraint generation iteration for each problem?
  3. Q3:
    How can we explain the results from the worst-case analysis to gain further insight about a policy?
To answer these questions, we evaluate on linear inventory control, linear reservoir management, and nonlinear VTOL control domains. Full descriptions are provided in the Appendix, including analysis of an additional nonlinear Boolean-action “intercept” domain. Inventory control has provably optimal PWS-S policies, whereas no optimal policy class is known explicitly for reservoir control; this allows us to answer Q1. The incorporation of the nonlinear VTOL problem will allow us to evaluate how CGPO scales across a variety of expressivity classes for linear/nonlinear dynamics. A public github repository [^1] Github URL suppressed for anonymous review. allows reproduction of all results and application of CGPO to arbitrary RDDL domains.
Domain Descriptions
@@ 3. Q3:
State $s_{t}$ describes the discrete level of stock available for a single good, action $a_{t}$ is the discrete reorder quantity, demand $\xi_{t}$ is stochastic and distributed as discrete uniform. Backlogging of inventories is allowed and is represented by negative $s_{t}$. The reward function consists of unit costs for product purchasing, excess and shortage of inventory. A planning horizon of $T=8$ is used. For this domain and reservoir below, we focus on learning factorized piecewise policies, i.e. C, S, PWS-C and PWS-S.
@@
The objective is to manage the water levels in a system of interconnected water reservoirs. State $s_{t,r}$ represents the continuous water level of reservoir $r$ with capacity $M_{r}$, action $a_{t,r}$ is the continuous amount of water released, and rainfall $\xi_{t,r}$ is a clipped normally-distributed random variable (to allow possibility of no rainfall). Each reservoir $r$ is connected to a set $U(r)$ of upstream reservoirs. Reward linearly penalizes any excess water level above $H_{r}$ and below $L_{r}$. We use $T=10$.
Nonlinear VTOL Control
Lastly, we consider a highly nonlinear aircraft VTOL control problem illustrated in Fig. 3. The objective is to balance two masses on opposing ends of a long pole, by applying a force to one of the two masses. The state consists of the angle $\theta_{t}$ and angular velocity $\omega_{t}$ of the pole, and the action is the force $F_{t}\in[0,1]$ applied to the mass. Time is discretized into intervals of length $\tau=0.1$ seconds, and we set $T=6$. The reward penalizes the difference between the pole angle and a target angle, as well as the magnitude of the force to ensure smoothness. We optimize for nonlinear quadratic policies.
Figure 3: The nonlinear VTOL control problem.
Nonlinear Intercept
To demonstrate an example of Discrete (Boolean) action policy optimization over nonlinear continuous state dynamics with independently moving but interacting projectiles, we turn to an intercept problem inspired by Scala et al. (2016). Space restrictions require us to relegate details and results to the Appendix, but we remark here that CGPO is able to derive an optimal policy.
Figure 4: Simulated return (top row) and worst-case error $\varepsilon^{*}$ (bottom row) on inventory (left column), reservoir (centre column) and VTOL (right column) over 100 independent roll-outs as a function of the number of iterations of constraint generation. Bars represent 95% confidence intervals calculated using 10 independent runs of Algorithm 1.
Empirical Results
MIPs are solved to an optimality gap of 5% and we use $p=0.995$. Fig. 2 illustrates examples of policies learned in a typical run of Algorithm 1. Fig. 4 compares the empirical (simulated) returns and error bounds $\varepsilon^{*}$ (optimal objective value of the inner problem) across different policy classes.
As hypothesized, CGPO can recover exact $(s,S)$ control policies for inventory control in the PWS-S class, which together with S policies achieve the lowest error and best return across all policy classes. Interestingly, as Fig. 4 shows, C and PWS-C policies perform relatively poorly even for $K=2$, but as expected, the error decreases with the policy expressivity class and the number of cases. This is intuitive, as Fig. 2 shows, the reorder quantity is strongly linearly dependent on the current stock, which is not easily represented as PWS-C. On the contrary, for reservoir control the class of PWS-C policies perform comparatively well for $K\geq 1$, and achieve (near)-optimality despite requiring a larger number of iterations. As expected, policies typically release more water as the level approaches the upper target bound. Finally, the optimal policy for VTOL applies large upward force when the angle or angular velocity are negative, with relative importance being placed on angle. As Fig. 2 shows, the force is generally greater when either the angle is below the target angle or the angular velocity is negative, and equilibrium is achieved by applying a modest upward force between the initial angle (0.1) and the target angle (0).
Inventory Reservoir
Policy Inner Outer Inner Outer
C 8 / 121 / 72 96 / 673 / 385 20 / 0 / 822 300 / 0 / 5823
PWS-S 32 / 129 / 72 384 / 772 / 387 80 / 0 / 842 1200 / 0 / 7133
C 160 / 0 / 64 780 / 0 / 480 540 / 0 / 320 4035 / 0 / 2400
PWS-S 168 / 0 / 112 709 / 168 / 1056 560 / 0 / 440 3797 / 540 / 4200
Table 2: Top half shows the number of binary/integer/real-valued decision variables in the inner and outer optimization problems of Algorithm 1 for different benchmark problems and policy classes at the last iteration. Bottom half shows the number of linear/quadratic/general constraints.
Figure 5: Worst-case state trajectory (left col.), actions (middle col.), and disturbances/noise (right col.) for C (top row) and S policies (bottom row) computed by Algorithm 1 for the reservoir control problem.
Worst-Case Analysis
Fig. 5 plots the state trajectory, actions and noise (i.e. rainfall) that lead to worst-case performance of C and S policies after the last iteration of Algorithm 1 for reservoir. Here, we observe that the worst-case performance for the optimal constant-value (C) policy occurs when the rainfall is low and both reservoirs become empty. Given that the cost of overflow exceeds underflow, this is expected as the C policy must release enough water to prevent high-cost overflow events during high rainfall at the risk of water shortages during droughts – this scenario results in non-zero cost and is identified by CGPO as incurring the worst-case cumulative reward in this policy class. In contrast, the optimal linear-valued (S) policy maintains safe water levels even in the worst-case scenario, since it can avoid underflow and overflow events by assigning release amounts that increase linearly with the water levels.
Analysis of Problem Size
To understand how the problem sizes in CGPO scale across differing expressivity classes of policies and dynamics, Table 2 shows the number of decision variables and constraints in the MIPs at the last iteration of constraint generation (more details are provided as supplementary material). For VTOL with a quadratic (Q) policy, the number of variables in the inner and outer problems are 0 / 0 / 212 and 0 / 0 / 512, resp., while the number of constraints are 114 / 18 / 72 and 270 / 95 / 155. The number of decision variables/constraints are fixed in the inner problem and grow linearly with iterations in the outer problem. Fortunately, the number of variables and constraints do not grow significantly with the expressiveness of the policy class.
Conclusion
We presented a novel bi-level mixed-integer programming formulation for DC-MDP policy optimization in predefined expressivity classes with bounded optimality guarantees. Constraint generation allowed us to decompose this into an inner problem that produces an adversarial worst-case constraint violation (i.e., trajectory), and an outer problem that improves the policy w.r.t. the most violated constraints. The result is our novel constraint generation policy optimization (CGPO) algorithm. Using inventory, reservoir and VTOL control problems, we demonstrated that the policies learned and their worst-case performance/failure modes were easy to interpret, while maintaining a manageable number of variables and constraints in the corresponding MIPs.
Appendix
Proof of Theorem 1
In the notation of Blankenship and Falk (1976), we define $X^{0}=[0,2BT]$, $X=\mathcal{W}$, $Y=\mathcal{A}^{T}\times\mathcal{S}_{1}\times\Xi_{p}^{T}$. Making the change of variable $\varepsilon=x_{0}\in X^{0}$, $x=\mathbf{w}$, and $y=(\alpha,\mathbf{s}_{1},\xi_{1:T})\in Y$, and defining the continuous function $f(x,y)=\varepsilon(\alpha,\mathbf{s}_{1},\xi_{1:T})=V(\alpha,\mathbf{s}_{1},\xi_{1:T})-V(\mathbf{w},\mathbf{s}_{1},\xi_{1:T})$, the problem (8) is equivalent to the problem:
$\min_{(x_{0},x)\in X^{0}\times X}$ $\,x_{0}$
$\mathrm{s.t.}$ $\,f(x,y)-x_{0}\leq 0,\quad\forall y\in Y$
as discussed on p. 262 of Blankenship and Falk (1976). Similarly, our Algorithm 1 is a special case of the Algorithm 2.1 in the aforementioned paper, specialized to the problem state, with $x_{t}=\mathbf{w}_{t}^{*}$. Thus, the assumption of Theorem 2.1 in Blankenship and Falk (1976) holds, and $\mathbf{w}_{t}^{*}$ is a solution of (8) as claimed. $\square$
Further Remarks About Theorem 1
In general, Algorithm 1 may not necessarily terminate, unless much stronger convexity requirements are placed on $V(\alpha,\mathbf{s})$ and $V(\mathbf{w},\mathbf{s})$, and potentially the class of MDPs solved. Thus, in a typical application, one can instead monitor the solution differences $\|\mathbf{w}_{t}^{*}-\mathbf{w}_{t-1}^{*}\|$, and terminate with an approximate solution when this difference is sufficiently small. However, our experiments run a fixed number of iterations, showing empirically that Algorithm 1 does indeed terminate on all problems with an exact optimal solution for most policy classes considered, at which point both the return and error curves “plateau” (cf. Figure 4).
Domains
Here we provide more complete mathematical descriptions of the benchmark problems used in the experiments. In what follows, we define $(\cdot)_{+}=\max[0,\cdot]$.
Linear Inventory
The state transition model can be written as:
$s_{t+1}$ $=s_{t}+a_{t}-\xi_{t},\quad\xi_{t}\sim\lfloor\mathrm{Uniform}(L,U)\rfloor,$
s.t. $0\leq a_{t}\leq B,\quad\forall t.$
We define costs $C,P$ and $S$ which represent, respectively, the purchase cost, excess inventory cost and shortage cost. The reward function can thus be written as
$r(\mathbf{s}_{t+1})$ $=-C\times a_{t}-P\times(s_{t+1})_{+}-S\times(-s_{t+1})_{+}.$
We set $B=10$, $C=0.5$, $P=S=2$, $L=2$, $U=6$. If $P>C$ and $T=\infty$, then a PWS-S policy is provably optimal (otherwise if $T<\infty$, then it may be non-stationary).
Linear Reservoir
Let $s_{t,r}$ denote the water level of reservoir $r$, $a_{t,r}$ be the amount of water to release from reservoir $r$, $\xi_{t,r}$ denote the stochastic rainfall amount affecting reservoir $r$, $U(r)$ denote the set of upstream reservoirs connected to reservoir $r$, and $M_{r}$ be the maximum capacity of reservoir $r$. The state transition model can be written as:
$s_{t+1,r}$ $=\min\Big{[}M_{r},\Big{(}s_{t,r}+(\xi_{t,r})_{+}-a_{t,r}+\sum_{d\in U(r)}a_{t,d}\Big{)}_{+}\Big{]},$
$\xi_{t,r}$ $\sim\mathrm{Normal}(m_{r},v_{r}),$
s.t. $0\leq a_{t,r}\leq s_{t,r},\quad\forall t,r.$
The goal is to maintain the level of water within target range $[L_{r},H_{r}]$. Penalties $l_{r}$ and $h_{r}$ are assigned for any quantity of water above or below the required target level. Thus, the reward function can be written as
@@ Conclusion
Our experiment uses a two-reservoir system with the following values for reservoir 1 $L_{1}=20,H_{1}=80,M_{1}=100,m_{1}=5,v_{1}=5,U(1)=\emptyset$, and the following values for reservoir 2 $L_{2}=30,H_{2}=180,M_{2}=200,m_{2}=10,v_{2}=10,U(2)=\{1\}$. Costs are shared and set to $l_{r}=-10,h_{r}=-100$.
Nonlinear VTOL Control
The state dynamics evolve according to:
@@
The reward is defined as the negative absolute difference between the angle at each epoch and the target angle $\theta_{target}=0$, i.e.
@@
We use values $l_{1}=1,l_{2}=0.5,m_{1}=10,m_{2}=1,h=0.4,g=9.8$.
Quadratic policies are of the form
$F(\theta_{t},\omega_{t})=w_{0}+w_{1}\theta_{t}+w_{2}\omega_{t}+w_{3}\theta_{t}\omega_{t}+w_{4}\theta_{t}^{2}+w_{5}\omega_{t}^{2},$
where $w_{i}$ are bounded in $[-100,100]$. Since coefficients are allowed to be both positive and negative, this increases the expressivity of the policy class at the expense of potentially inducing non-convexity of the objective.
@@
The intercept problem is a nonlinear domain with Boolean actions. It involves two objects moving in continuous trajectories in a subset of $\mathbb{R}^{2}$, in which one object (e.g. missile, bird) flies in a parabolic arc across space towards the ground, and must be intercepted by a second object (e.g. anti-ballistic missile, predator) that is fired from a fixed position on the ground. While the problem is best described in continuous time, we study a discrete-time version of the problem. The state includes the position $(x_{t},y_{t})$ of the missile at each decision epoch, as well as “work” variables indicating whether the interceptor has already been fired $f_{t}$ as well as its vertical elevation $i_{t}$. Meanwhile, the action $a_{t}$ is Boolean-valued and indicates whether the interceptor is fired at a given time instant.
The state transition model can be written as:
$x_{t+1}$ $=x_{t}+v_{x},\quad y_{t+1}=h-\frac{1}{2}gx_{t}^{2}$
$f_{t+1}$ $=f_{t}\lor a_{t},\quad i_{t+1}=\begin{cases}i_{t}&\mbox{ if }f_{t}=0\\ i_{t}+v_{y}&\mbox{ if }f_{t}=1\end{cases},$
$r(s_{t+1})=\begin{cases}1&\mbox{ if }f_{t+1}\land|x_{t+1}-i_{x}|\leq\delta\land|y_{t+1}-i_{t+1}|\leq\delta\\ 0&\mbox{ otherwise}\end{cases}.$
We set $v_{x}=0.1,h=5,g=9.8,i_{x}=0.7$.
where $a_{1},a_{2}\in\{0,1\}$ and $l,u,w_{1},w_{2}$ are all tunable parameters.
Domain RDDL Descriptions
RDDL domain descriptions of the above problems are provided as accompaniment to the supplementary material.
@@
For reservoir and inventory control, bounds on the initial state have been chosen as intervals centred on the initial state specified in the RDDL domain description. For reservoir control, these bounds are intervals centred around the initial state $(75,150)$, i.e. $[50,100]$ for reservoir 1 and $[100,200]$ for reservoir 2. For inventory control, the bounds are $[0,2]$, for VTOL they are fixed to the initial state of the system $\theta=0.1,\omega=0$, and for intercept they are also fixed to the initial state $x=0,y=5,f=0,i=0$.
Gurobi and Environment Settings
We use the Python implementation of the Gurobi Optimizer (GurobiPy) version 10.0.1, build v10.0.1rc0 for Windows 64-bit systems, with an academic license. Default optimizer settings have been used, with the exception of NumericFocus set to 2 in order to enforce numerical stability, and the MIPGap parameter set to 0.05 to terminate when the optimality gap reaches 5%. All experiments were conducted on a single machine with an Intel 10875 processor (base at 2.3 GHz, overclocked at 5.1 GHz) and 32 GB of RAM.
Additional Implementation Details
Action Constraints
Handling Numerical Errors
$y==1$ $\implies a\geq b+\epsilon$
$y==0$ $\implies a\leq b.$
Additional Experimental Results
Intercept
Fig. 6 illustrates the return and worst-case error of Boolean-valued policies for the intercept problem. Only the policies with at least one case are optimal, with corresponding error of zero. As illustrated in the last plot in Fig. 7, the optimal policies fire whenever a threat is detected in the top right corner of the airspace.
Policy Representations
@@
Fig. 8 and Fig. 9 showcase policies learned in other expressivity classes besides those shown in the main text, for inventory and reservoir control, respectively. Fig. 7 illustrates the policies for the intercept problem. As can be seen, policies generally remain intuitive and easy to explain even as the complexity increases. For instance, PWS2-C policies for inventory control order more inventory as the stock level decreases, while the amount of water released is increasing in the water level for reservoir control. Interestingly, the PWS2-C policies for inventory control and the intercept problem contain redundant constraints.
Worst-Case Analysis
Fig. 10 and Fig. 11 illustrate worst case scenarios induced by the C and S policies for inventory and VTOL control, respectively. Similar to reservoir control, a non-trivial worst-case scenario is identified by CGPO for inventory control for optimal C policies but not S policies. This scenario corresponds to very high demand that exceeds the constant reorder quantity, causing stock-outs. For the VTOL control problem, there is no worst-case scenario with non-zero cost upon convergence at the final iteration of CGPO (bottom row). We have also shown the worst-case scenario computed at the first iteration prior to achieving policy optimality, where a worst-case scenario corresponds to no force being applied causing the system to eventually lose balance. The ability to interpret the worst-case scenario prior to convergence, or if convergence cannot be achieved, could help to decide whether the policy is trustworthy, and how to prepare for the worst-case if the policy is deployed in practice.
Analysis of Problem Size
Fig. 12, Fig. 13 and Fig. 14 summarize the total number of variables and constraints in the outer MIP formulations solved by CGPO at each iteration for inventory, reservoir and VTOL control, respectively. Each iteration of constraint generation requires a roll-out from the dynamics and reward model, which in turn requires instantiate a new set of variables to hold the intermediate state and reward computations, and thus the number of variables and constraints grows linearly with the iterations. Even at the last iteration of CGPO, we see that the number of variables and constraints remains manageable.
Figure 6: Simulated return (left) and optimal value of $\varepsilon^{*}$ (right) as a function of the number of iterations of constraint generation for the intercept domain.
$fire=1$
Figure 7: From left to right, examples of learned B, PWS1-B and PWS2-B policies and visualization of PWS-B policy for the intercept problem.
@@
Figure 8: From left to right, examples of learned C, S, PWS1-C and PWS2-C policies for the inventory control problem.
@@
Figure 9: From left to right, examples of learned C, S, PWS2-C and PS1-S policies for the reservoir control problem.

### v2

@@ 3.4 Problem Expressivity Class Analysis
Table 1: Problem classes of the inner/outer optimization problems in CGPO for different expressivity classes of dynamics/reward (columns) and policies (rows).
In Table 1, we present the relationship between the expressivity classes of policies and state transition dynamics and the corresponding classes of the inner and outer optimization problems. The policy classes of interest include PWS-C, PWL-C, PWS-L, PWL-L, PWN-N and the different variants of piecewise polynomial policies. Meanwhile, expressivity classes for state dynamics and reward include linear, polynomial and general nonlinear functions (their piecewise counterparts generally fall under the same expressivity classes, and are excluded for brevity). Interestingly, as shown in Appendix 6.2, when the policy and state dynamics are both linear, the outer problem is a PP. In a similar vein, a PWL-C policy and linear dynamics result in a MIBCP outer problem due to the bilinear interaction between successor state decision variables and policy weights in the linear conditions.
Our experiments in the next section empirically evaluate PWS-C and PWS-L policies with linear dynamics and Q policies with nonlinear dynamics. This requires solving mixed-integer problems with large numbers of decision variables ranging from MILP to PP to MINLP.
4 Empirical Evaluation
We empirically validate the performance of CGPO on several MDPs, aiming to answer the following questions:
  1. Q1
@@ 4 Empirical Evaluation
  2. Q2
    How do different policy expressivity classes perform for each problem?
  3. Q3
    Does the worst-case analysis provide further insight about a policy?
4.1 Domain Descriptions
To answer these questions, we evaluate on linear Inventory, linear Reservoir, and nonlinear VTOL (vertical take-off and landing) domains as summarized below. Inventory has provably optimal PWS-S policies, whereas no optimal policy class is known explicitly for Reservoir. Experimental details and additional results are discussed in Appendix 6.4 and 6.5 respectively, and a public GitHub repository [^2] https://github.com/pyrddlgym-project/pyRDDLGym/tree/GurobiCompilerBranch/ allows reproduction of all results and application of CGPO to arbitrary domains.
@@ 4.1 Domain Descriptions
State $s_{t}$ describes the discrete level of stock available for a single good, action $a_{t}\in[0,B]$ is the discrete reorder quantity, and demand $\xi_{t}$ is stochastic and distributed as discrete uniform:
$s_{t+1}$ $=s_{t}+a_{t}-\lfloor\xi_{t}\rfloor,\quad\xi_{t}\sim\mathrm{Uniform}(L,U).$
We permit backlogging of inventories, represented as negative $s_{t}$. We define costs $C,P$ and $S$ which represent, respectively, the purchase cost, excess inventory cost and shortage cost, and the reward function can be written as
$r(\mathbf{s}_{t+1})$ $=-C\times a_{t}-P\times(s_{t+1})_{+}-S\times(-s_{t+1})_{+},$
where $(\cdot)_{+}=\max[0,\cdot]$. We set $B=10$, $C=0.5$, $P=S=2$, $L=2$, $U=6$. If $P>C$ and $T=\infty$, then a PWS-S policy is provably optimal (otherwise if $T<\infty$, then it may be non-stationary). A planning horizon of $T=8$ is used. For this domain and reservoir below, we focus on learning factorized piecewise policies, i.e. C, S, PWS-C and PWS-S.
@@
The goal is to manage the water level in a system of interconnected reservoirs. State $s_{t,r}$ represents the continuous water level of reservoir $r$ with capacity $M_{r}$, action $a_{t,r}\in[0,s_{t,r}]$ is the continuous amount of water released, and rainfall $\xi_{t,r}$ is a truncated normally-distributed random variable. Each reservoir $r$ is connected to a set $U(r)$ of upstream reservoirs, thus:
$s_{t+1,r}=\min\Big{[}M_{r},\Big{(}s_{t,r}+(\xi_{t,r})_{+}-a_{t,r}+\sum_{d\in U(r)}a_{t,d}\Big{)}_{+}\Big{]},\quad\xi_{t,r}\sim\mathcal{N}(m_{r},v_{r}).$
Reward linearly penalizes any excess water level above $H_{r}$ and below $L_{r}$:
@@
Our experiment uses a two-reservoir system with the following values for reservoir 1 $L_{1}=20,H_{1}=80,M_{1}=100,m_{1}=5,v_{1}=5,U(1)=\emptyset$, and the following values for reservoir 2 $L_{2}=30,H_{2}=180,M_{2}=200,m_{2}=10,v_{2}=10,U(2)=\{1\}$. Costs are identical for all reservoirs and are set to $l_{r}=-10,h_{r}=-100$. We use $T=10$.
Nonlinear VTOL
The goal is to balance two masses on opposing ends of a long pole. The state consists of the angle $\theta_{t}$ and angular velocity $\omega_{t}$ of the pole, and the action is the force $F_{t}\in[0,1]$ applied to the mass:
Figure 2: Nonlinear VTOL system.
@@
Time is discretized into intervals of $\tau=0.1$ seconds, and we set $T=6$. The reward penalizes the difference between the pole angle and a target angle:
@@
We use values $l_{1}=1,l_{2}=0.5,m_{1}=10,m_{2}=1,h=0.4,g=9.8$. We optimize for nonlinear quadratic policies.
@@
To demonstrate an example of discrete (Boolean) action policy optimization over nonlinear continuous state dynamics with independently moving but interacting projectiles, we turn to the Intercept problem (Scala et al., 2016). Space restrictions require us to relegate details and results to Appendix 6.3 and 6.5, but we remark here that CGPO is able to derive an optimal policy.
@@
For Reservoir, the initial state bounds are set to $[50,100]$ for reservoir 1 and $[100,200]$ for reservoir 2. For Inventory, the bounds are $[0,2]$, and for VTOL and Intercept they are fixed to the initial state of the system.
Figure 3: Reservoir (left), Inventory (middle), VTOL (right): Simulated return (top row) and worst-case error $\varepsilon^{*}$ (bottom row) over 100 roll-outs as a function of the number of iterations of constraint generation. Bars represent 95% confidence intervals calculated using 10 independent runs of CGPO.
$F=0.6-15.4\theta-2.3\omega+100\theta^{2}-1.5\omega^{2}$
Figure 4: Reservoir (left), Inventory (middle), VTOL (right): Examples of optimal policies computed at the end of CGPO.
4.2 Empirical Results
Fig. 3 compares the empirical (simulated) returns and error bounds $\varepsilon^{*}$ (optimal objective value of the inner problem) across different policy classes. Fig. 4 illustrates examples of policies learned in a typical run of CGPO. As hypothesized, CGPO can recover exact $(s,S)$ control policies for Inventory in the PWS-S class, which together with S policies achieve the lowest error and best return across all policy classes. Interestingly, as Fig. 3 shows, C and PWS-C policies perform relatively poorly even for $K=2$, but as expected, the error decreases as the expressivity of the policy class, and in particular the number of cases, increases. This is intuitive, as Fig. 4 shows, the reorder quantity is strongly linearly dependent on the current stock, which is not easily represented as PWS-C. In constrast, on Reservoir the class of PWS-C policies perform comparatively well for $K\geq 1$, and achieve (near)-optimality despite requiring a larger number of iterations. As expected, policies typically release more water as the water level approaches the upper target bound. Finally, the optimal policy for VTOL applies a large upward force when the angle or angular velocity are negative, with relative importance being placed on angle. As Fig. 4 (right) shows, the force is generally greater when either the angle is below the target angle or the angular velocity is negative, and equilibrium is achieved by applying a modest upward force between the initial angle (0.1) and the target angle (0).
@@ 4.2 Empirical Results
Fig. 5 and 6 provide examples of policies learned in other expressivity classes for Inventory and Reservoir problems, respectively. In all cases, policies remain intuitive and easy to explain even as the complexity increases. For instance, PWS2-C policies for Inventory order more inventory as the stock level decreases, while for Reservoir, the amount of water released is increasing in the water level.
@@
Figure 5: Inventory: Examples of C, S, PWS1-C and PWS2-C (left to right) policies computed by CGPO.
@@
Figure 6: Reservoir: Examples of C, S, PWS2-C and PS1-S (left to right, top to bottom) policies computed by CGPO.
4.3 Ablations